OpenStudy (waheguru):

i dont understan this please help a line is perpindcular to 4x+7y-1=0 and has the same x-intersept as 2x-5y=8

OpenStudy (waheguru):

FIND THE EQUATION of the line

OpenStudy (akshay_budhkar):

do you know the equation of the line is y=mx+c where m is the slope and c is the intercept

OpenStudy (slaaibak):

It means that they (the line you are looking for that is perpendicular to 4x + 7y -1=0, and the 2x-5y=8) intercept each other at y=0

OpenStudy (waheguru):

yes i know that all but what is the equaiton of that line above

OpenStudy (slaaibak):

do you know what it means when two lines are perpendicular? m1 * m2 = -1

OpenStudy (slaaibak):

Use that to solve. We are here to assist, not to do it for you :)

OpenStudy (waheguru):

nice answer VERRRRY HELPFUL

OpenStudy (slaaibak):

Pleasure, always here to help!